Find the values of x given X2 - 4x + 13 = 0


1
Expert's answer
2021-02-24T06:00:35-0500

The roots of qudratic equation are:


"x_{1,2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a},"


where a = 1;

b = -4;

c = 13.


So,


"x_{1,2}=\frac{-(-4)\pm \sqrt{(-4)^2-4\cdot 1\cdot13}}{2\cdot 1}=2\pm 3\sqrt{-1}=2\pm 3i."


Answer: x1 = 2 + 3i, x2 = 2 - 3i.